# Mask Image Watermarking

### v1

@@ 3.8 Usage Scenarios
The following section outlines the usage scenarios for the two variants, MaskMark-D and MaskMark-ED, highlighting their distinct watermark embedding strategies tailored to different protection needs. While MaskMark-D is designed for comprehensive image protection, MaskMark-ED focuses on safeguarding specific regions of interest.
MaskMark-D.
MaskMark-D operates without a mask in the encoder, supporting global watermark embedding and local watermark extraction. The mask is introduced only during decoder training, enabling selective focus on watermarked regions for accurate extraction. This makes MaskMark-D ideal for full-image protection. Even in cases of heavy tampering, where only part of the watermark remains, MaskMark-D can still extract the watermark and identify tampered areas. Its robustness under such conditions makes it a reliable solution for comprehensive image protection.
MaskMark-ED.
In contrast, MaskMark-ED uses a mask in the encoder, allowing for localized watermark embedding and extraction. This approach is best suited for protecting specific, valuable regions of an image, rather than the entire image. The focus is not on tamper localization across the whole image but on ensuring that key content can be traced if extracted and misused. This targeted protection makes MaskMark-ED ideal for scenarios where only particular regions need safeguarding against unauthorized use.
@@ 4.1 Implementation Details
For all experiments, we train MaskMark on 83k images from the MS-COCO (Lin et al., 2014) 2014 training set. All images are resized and center-cropped to $256\times 256$. Training is conducted for 100k steps with a batch size of 16 on a single NVIDIA A6000 GPU. We use the AdamW optimizer with a learning rate of $1\times 10^{-4}$, and apply a cosine learning rate scheduler with 2k warm-up steps. We adopt an easy-to-hard training strategy inspired by TrustMark (Bui et al., 2023). During the first 0.5k steps, the mask is set to full (i.e., all ones) and no distortion is applied. From step 0.5k to 1k, we introduce all types of masks. After 1k steps, distortions are added. The encoder loss weight $\beta_{enc}$ is fixed at 1, while the decoder loss weight $\beta_{dec}$ is initially set to 20 and linearly decayed to 0.2 over the first 5k steps. The mask loss weight $\alpha$ is set to 0.5. The JND module in the encoder is introduced and tuned starting from step 5k, with the scaling factor $\mu$ set to 1. The details of the distortion pool used during training are provided in Appendix A.1.
For evaluation, we fix the image resolution to $512\times 512$. If a baseline method does not support this resolution, we adopt the resolution scaling strategy from TrustMark (Bui et al., 2023) to interpolate the watermark strength, which preserves watermarking performance. To ensure comparable visual fidelity across variants, we set the JND scaling factor $\mu$ for MaskMark-D and MaskMark-ED to 1.3 and 1.75, respectively. For robustness evaluation, we separately assess valuemetric and geometric distortions. For valuemetric robustness, we randomly sample from a set of ten common distortions, including JPEG Compression, Gaussian Filter, Gaussian Noise, Median Filter, Salt&Pepper Noise, Resize, Brightness, Contrast, Hue, and Saturation. For geometric robustness, we randomly sample from three typical transformations: Rotation, Perspective, and Horizontal Flip. These distortions collectively cover the vast majority of real-world transformations that watermarked images are likely to encounter in practical scenarios. Detailed parameter settings for each distortion are provided in Appendix A.2.
@@ 4.2 Global and Local Watermarking Comparison
We compare MaskMark against seven recent open-source watermarking methods with state-of-the-art performance. Among them, StegaStamp (Tancik et al., 2020), SepMark (Wu et al., 2023), TrustMark (Bui et al., 2023), EditGuard (Zhang et al., 2024b), Robust-Wide (Hu et al., 2024b), and VINE (Lu et al., 2024) are designed for global watermarking. WAM (Sander et al., 2025), on the other hand, is a dedicated method for local watermarking. For EditGuard, we evaluate two publicly available variants: the clean version (EditGuard-C), trained without noise augmentation, and the degraded version (EditGuard-D), trained with noise. For VINE, we use the more robust variant, VINE-R.
For global watermarking evaluation, we randomly select 1k images from the MS-COCO 2014 validation set. We use PSNR and SSIM to assess the visual quality of the watermarked images, and Bit Accuracy to evaluate the watermark extraction performance. For local watermarking evaluation, we split the 41k images in the MS-COCO 2014 validation set into 12 subsets based on the ratio of the masked area to the entire image: 1–5%, 5–10%, 10–20%, …, 80–90%, 90–95%, and 95–99%. From each subset, we randomly select 400 images and then invert the masked regions to simulate both inpainting and outpainting scenarios. That is, if the original masked area is $a$$b$%, the inverted masked area will cover $(100{-}b)$$(100{-}a)$of the image. This results in $400\times 2=800$ images per subset, yielding a total of $800\times 12=9600$ image-mask pairs for evaluation. By default, MaskMark is configured to embed 32 bits for fair comparison with WAM. This capacity is generally sufficient for local watermark embedding and extraction. However, unlike WAM, MaskMark is not restricted to 32-bit capacity. We provide additional results for varying bit lengths in Sec. 4.5.
For the baseline methods and our MaskMark-D, we embed watermark bits across the entire image and then replace the unmasked regions with the corresponding original pixels to obtain the final local watermarked image. In contrast, our MaskMark-ED leverages the mask to embed watermark bits only within the masked region, and similarly replaces the unmasked areas with original pixels, ensuring that the resulting image contains localized watermarking restricted to the target area.
@@
The global watermarking results for all methods are summarized in Table 1. First, both MaskMark-D and MaskMark-ED achieve high visual fidelity, with PSNR scores above 39.5 and SSIM scores exceeding 0.98. These results outperform WAM and are only marginally lower than TrustMark and Robust-Wide. More importantly, under this high-fidelity setting, both variants of MaskMark still maintain near 100% bit accuracy, even under various valuemetric and geometric distortions. This demonstrates significantly better robustness compared to both global and local watermarking baselines. Notably, geometric distortions pose a considerable challenge for existing global watermarking methods, often leading to failure. In contrast, our MaskMark framework remains robust and reliable in such scenarios.
Table 1: Comparison with other baseline watermarking methods in terms of global watermarking. The best and the second best results are highlighted in bold and underlined, respectively.
@@
StegaStamp 100 28.87 0.9019 0.9990 0.9976 0.6646
SepMark 30 35.73 0.9876 0.9957 0.9643 0.5086
TrustMark 100 41.19 0.9922 0.9996 0.9955 0.7868
EditGuard-C 64 37.27 0.9332 0.9991 0.5482 0.4925
EditGuard-D 64 32.30 0.8199 0.9999 0.5444 0.4975
Robust-wide 64 41.58 0.9923 1.0000 0.9944 0.4951
VINE 100 36.04 0.9874 0.9997 0.9986 0.5012
@@
WAM 32 39.32 0.9791 1.0000 0.9986 0.8979
MaskMark-D 32 39.55 0.9814 1.0000 1.0000 0.9998
MaskMark-ED 32 39.52 0.9828 1.0000 1.0000 1.0000
@@
The local watermarking results for all methods are presented in Figure 2. First, we observe that the extraction accuracy of global watermarking methods drops significantly as the watermarked region becomes smaller. This highlights their limitation in handling local watermarking tasks, especially when only a small portion of the image contains the watermark. In contrast, both WAM and our proposed MaskMark maintain high extraction accuracy even when the watermarked area is limited. Among them, MaskMark consistently outperforms WAM, with a more pronounced advantage under distortion scenarios, demonstrating its superior robustness and reliability in practical settings. Finally, comparing our two variants, MaskMark-ED performs better than MaskMark-D when the watermark region is small. As the watermarked area increases, the performance gap between the two variants narrows.
Figure 2: Watermark extraction performance of different methods under different ratios of watermarked area. The intervals of ratios are: 1-5%, 5-10%, …, 95-99%, 99-100%. We select the average value for each interval’s ratios to stand for the interval (e.g., 3% for 1-5%).
The visualized watermark patterns embedded by MaskMark-D and MaskMark-ED are shown in Figure 3, with additional results provided in Appendix B.1.
Figure 3: Visualization results of global watermark embedding using MaskMark-D and local watermark embedding using MaskMark-ED.
@@ 4.3 Watermark Localization Comparison
We compare MaskMark with EditGuard (Zhang et al., 2024b) and WAM (Sander et al., 2025), two methods with watermark localization capabilities, selected from our earlier baseline comparisons in Sec. 4.2. Localization performance is evaluated using the local watermarking dataset described in Sec. 4.2, with Intersection-over-Union (IoU) metrics computed between the predicted and ground-truth regions for both the watermarked and unwatermarked areas.
@@ Results.
The localization results of different methods are presented in Figure 4. First, our method MaskMark consistently achieves the best localization performance across nearly all watermark ratios and distortion conditions, with WAM showing a noticeable performance gap and EditGuard performing significantly worse. Second, MaskMark-ED achieves better localization accuracy in watermarked regions than MaskMark-D when the watermarked area is small, while MaskMark-D demonstrates superior performance in localizing unwatermarked regions when the watermarked area is large. This trend becomes more pronounced under distortion conditions. An interesting observation is that EditGuard-C exhibits an abnormal increase in localization accuracy for unwatermarked regions as the watermarked area becomes larger, a behavior not seen in EditGuard-D. We hypothesize that this may be due to overfitting: since EditGuard-C is trained solely on clean images without distortions, it may become overly sensitive to specific patterns present in undistorted data. The visualized localization results of different methods are shown in Figure 5, with additional results provided in Appendix B.2.
Figure 4: Localization performance of different methods under different ratios of watermarked area.
Figure 5: Visualization results of watermark localization using different methods. The inpainting and outpainting results are obtained by applying stable-diffusion-2-inpainting (Rombach et al., 2022) to the masked regions for content reconstruction.
4.4 Multiple Watermarks Embedding Comparison
@@
We compare the performance of our MaskMark-ED with WAM (Sander et al., 2025) in a multi-watermark embedding setting. Following WAM’s setup, we embed different watermarks into 1–5 non-overlapping square-masked regions to evaluate performance as the number of embedded watermarks increases. These masks are placed at the center, top-left, top-right, bottom-left, and bottom-right positions, forming a checkerboard-like pattern when all five regions are used (see Appendix B.3). Unlike WAM, which assigns 10% of the image area to each mask, we limit each mask to just 5%, making the task significantly more challenging. For each watermark count from 1 to 5, we randomly sample 400 images from the MS-COCO 2014 validation set for evaluation. Performance is assessed using the average watermark extraction accuracy across multiple embedded watermarks and the mean IoU of the predicted watermark regions. For our MaskMark-ED model, we apply OpenCV’s cv2.connectedComponents function to segment the extracted mask into disjoint regions, allowing independent watermark extraction from each region.
@@ Results.
Figure 6 presents the comparison results. First, our MaskMark-ED consistently outperforms WAM in both watermark extraction and localization, across all tested numbers of embedded watermarks and under various distortion conditions. Second, despite being trained solely with single-watermark supervision, MaskMark-ED generalizes well to multi-watermark settings, demonstrating strong scalability. Under geometric distortions, we observe a degradation in extraction accuracy as the number of embedded watermarks increases. This is primarily due to the spatial transformations disproportionately shrinking the watermark regions located in the image corners, thereby reducing the effective area available for extraction (see Appendix B.3).
@@ 4.5 More Analysis
Unlike WAM (Sander et al., 2025), our MaskMark can be readily scaled to support longer watermark bit lengths. As shown in Figure 7, MaskMark maintains high extraction accuracy across various bit lengths. Even at 64 bits, the accuracy experiences only a slight drop and still outperforms WAM at 32 bits. While the accuracy decline becomes more noticeable at 128 bits, MaskMark consistently surpasses WAM under both no distortion and valuemetric distortion conditions. The only exception arises under geometric distortion, where a performance gap emerges when the watermark region covers between 5% and 75%. These results underscore the scalability and optimization-friendly design of our method, emphasizing its practical advantages.
Figure 7: Watermark extraction performance of MaskMark-D and MaskMark-ED with different bits length. We also put the results of WAM for comparison.
Effects of Different Visual Quality Enhancement Methods.
Our experiments demonstrate that using only an MSE loss in pixel space to constrain the difference between the watermarked and original images (referred to as the Base version) already yields high PSNR and SSIM scores. However, despite the favorable metrics, the resulting images often contain visible artifacts that may compromise perceived visual quality. To address this issue, we investigate several enhancement strategies: (1) incorporating a GAN (Isola et al., 2017) loss, (2) adding a perceptual constraint via LPIPS (Zhang et al., 2018) loss in the feature space, and (3) modulating the watermark signal using a Just Noticeable Difference (JND) (Wu et al., 2017) module. We conduct experiments on the MaskMark-D variant, but the findings similarly hold for MaskMark-ED.
As illustrated in Figure 8, when the weights of the GAN and LPIPS losses are small, they are insufficient to suppress artifacts effectively. Increasing these weights helps reduce artifacts but negatively impacts the performance of local watermark extraction and localization (see Figure 18 in Appendix B.4 for detailed results). In contrast, applying JND-based modulation proves more effective: it significantly reduces visible artifacts while maintaining performance comparable to the Base version. These findings suggest that, unlike global perceptual losses such as GAN or LPIPS, JND offers a more adaptive and content-aware modulation strategy. It effectively suppresses visual artifacts while preserving the watermark’s integrity, making it a practical choice for visual quality enhancement.
Figure 8: Visualization results of watermark using different visual quality enhancement methods. The residual image is acquired by $10\times|I_{wm}-I_{orig}|$ for observation. Zoom in to see more details. The numbers in parentheses indicate the corresponding loss weights. For example, GAN (0.001) means the GAN loss is assigned a weight of 0.001 in the total loss function, i.e., $\mathcal{L}_{\text{total}}=\beta_{enc}\mathcal{L}_{\text{enc}}+\beta_{dec}\mathcal{L}_{\text{dec}}+0.001\mathcal{L}_{\text{GAN}}$, and LPIPS (0.01) indicates that the LPIPS loss is weighted by 0.01, i.e., $\mathcal{L}_{\text{enc}}=\mathcal{L}_{\text{MSE}}(I_{\text{wm}},I_{\text{orig}})+0.01\mathcal{L}_{\text{LPIPS}}$. To ensure a fair comparison, we adjust either the strength of the added watermark residual or the JND modulation coefficient to maintain comparable PSNR and SSIM across different settings.
@@
Although MaskMark is trained with a broad set of common valuemetric and geometric distortions, it is impractical to account for all possible types of distortions during training. Fortunately, the flexibility of our framework allows users to easily perform task-specific enhancement and fine-tuning based on their particular needs. As a demonstration, we target a recently popular class of VAE-based adaptive attacks (Kingma & Welling, 2014), which use variational autoencoders to reconstruct images and potentially erase embedded watermark signals. Specifically, we expand the distortion pool during fine-tuning by incorporating VAE modules from Stable Diffusion v1-4 (Rombach et al., 2022), Bmshj18 (Ballé et al., 2018), and Cheng20 (Cheng et al., 2020). These VAE-based distortions are applied with a probability of 50%, while the original distortion types are retained with the remaining 50%. The hyperparameters are configured as follows: $\beta_{\text{dec}}=0.3$, learning rate $=1\times 10^{-4}$, and the quality levels for Bmshj18 and Cheng20 are set to 5.
We find that fine-tuning for 20k steps on a single A6000 GPU, approximately 5 hours, is sufficient to achieve strong performance. As shown in Figure 9, the fine-tuned MaskMark exhibits significantly enhanced robustness against VAE-based attacks, clearly outperforming existing baselines. However, we also observe a slight drop in robustness against other types of distortions. Despite this trade-off, MaskMark still maintains superior performance compared to existing methods overall. This highlights the adaptability of our approach, enabling users to make informed trade-offs based on their specific application requirements.
Figure 9: The effect of VAE fine-tuning on the robustness of our MaskMark. Fine-tuning the VAE enhances robustness against VAE attacks, with minimal impact on the original robustness performance.
@@
We evaluate watermark extraction performance using three types of masks during decoding: a full mask, a predicted mask, and a ground-truth mask. With the full mask, the entire image is passed to the decoder for extraction. The predicted mask retains only the regions identified as containing the watermark, setting the pixel values of all other regions to zero. The ground-truth mask isolates the actual watermark-containing regions based on annotation.
As shown in Figure 10, when the watermark is embedded in small, localized regions, using the full mask leads to significantly lower extraction accuracy. In contrast, the predicted mask yields a notable improvement, and the ground-truth mask further boosts performance. These results indicate that removing irrelevant image content helps reduce interference, thereby enhancing local watermark extraction. They also underscore the importance of performing localization prior to extraction in scenarios involving spatially localized watermarks.
@@
We compare MaskMark with EditGuard and WAM, both of which support watermark extraction and localization, in terms of training and inference costs. The results are summarized in Table 2. First, in terms of model size, MaskMark has fewer parameters than WAM, with the encoder and decoder having comparable sizes. In contrast, WAM allocates most of its parameters to the decoder, as its decoding task is inherently more complex. Second, regarding training overhead, MaskMark is significantly more efficient than WAM. Measured in TFLOPs, the training cost of MaskMark is approximately 1/15 that of WAM. Although EditGuard does not report explicit training overhead, it is noteworthy that even without incorporating strong distortions for robustness during training, its total data consumption (steps×$\times$×batch) already amounts to 60% of ours. Finally, for inference overhead, despite having fewer parameters, EditGuard embeds both localization and copyright watermarks into the image, which substantially increases both memory usage and inference time. In contrast, MaskMark and WAM incur much lower inference latency and exhibit similar memory consumption.
@@
EditGuard 3.6 2.6 6.2 25w 4 - - - - 0.074 s 0.080 s 0.154 s 2.15 GB
WAM 1.1 96.0 97.1 368w 16 8 V100 90 h $1.13\times 10^{16}$ 0.015 s 0.017 s 0.032 s 2.33 GB
MaskMark 31.1 32.2 63.3 10w 16 1 A6000 20 h $7.74\times 10^{14}$ 25.84 GB 0.009 s 0.022 s 0.031 s 2.21 GB

### v2

@@ 3.4 Usage Scenarios
This section outlines the usage scenarios of MaskMark-D and MaskMark-ED, which are tailored to different protection requirements. MaskMark-D is designed for full-image protection. It enables reliable watermark extraction even when parts of the image are tampered with, making it suitable for scenarios that demand global content integrity and tamper detection, such as copyright enforcement or image provenance verification. MaskMark-ED, on the other hand, focuses on protecting specific regions of interest, such as faces, logos, or sensitive content. It allows targeted verification or tracing if these regions are reused or misappropriated, without introducing watermarks across the entire image. This strategy not only supports privacy-aware or content-specific protection, but also tends to provide better robustness within the marked regions.
@@ 4.1 Implementation Details
Training.
For all experiments, we train MaskMark on 83k images from the MS-COCO 2014 training set [12] and the training details are provided in Appendix C.1.1.
Evaluation.
To ensure fair comparison, we fix the image resolution to $512\times 512$. For baseline methods that do not support this resolution, we follow the resolution scaling strategy from TrustMark [2] to interpolate the watermark strength (see Appendix C.2.1), which has been shown to preserve watermarking performance. To ensure comparable visual fidelity across variants, we set the JND scaling factor $\mu$ to 1.3 for MaskMark-D and 1.75 for MaskMark-ED. For robustness evaluation, we separately assess valuemetric and geometric distortions. For valuemetric robustness, we randomly sample from a set of ten common distortions, including JPEG Compression, Gaussian Filter, Gaussian Noise, Median Filter, Salt&Pepper Noise, Resize, Brightness, Contrast, Hue, and Saturation. For geometric robustness, we randomly sample from three typical transformations: Rotation, Perspective, and Horizontal Flip. These distortions collectively cover the vast majority of real-world transformations that watermarked images are likely to encounter in practical scenarios. Detailed parameter settings for each distortion are provided in Appendix C.2.2. We also evaluate robustness under each specific distortion individually and the detailed results are presented in Appendix D.2.
@@ 4.2 Global and Local Watermarking Comparison
We compare MaskMark with seven recent open-source watermarking methods, including both global (e.g., StegaStamp [21], SepMark [23], TrustMark [2], EditGuard [28], Robust-Wide [7], VINE [13]) and local (WAM [19]) approaches. We use the clean (EditGuard-C) and degraded (EditGuard-D) variants of EditGuard, and the robust version (VINE-R) of VINE. For global watermarking, we evaluate on 1k images from the MS-COCO 2014 validation set using PSNR, SSIM, and Bit Accuracy, where PSNR and SSIM measure the visual quality of watermarked images, and Bit Accuracy evaluates watermark extraction performance. For local watermarking, we use all 41k validation images and evaluate Bit Accuracy under different watermarked area ratios. By default, MaskMark embeds 32 bits for fair comparison with WAM, though MaskMark supports flexible bit lengths (see Sec. 4.5). See Appendix C.2.3 for more implementation details and evaluation settings.
@@
The global watermarking results for all methods are summarized in Table 1. First, both MaskMark-D and MaskMark-ED achieve high visual fidelity, with PSNR scores above 39.5 and SSIM scores exceeding 0.98. These results outperform WAM and are only marginally lower than TrustMark and Robust-Wide. More importantly, under this high-fidelity setting, both variants of MaskMark still maintain near 100% bit accuracy, even under various valuemetric and geometric distortions. This demonstrates significantly better robustness compared to both global and local watermarking baselines. Notably, geometric distortions, which often break existing global watermarking methods, are effectively handled by our robust and reliable MaskMark framework.
Table 1: Comparison with baseline watermarking methods in terms of global watermarking. The best and the second best results are highlighted in bold and underlined, respectively.
@@
StegaStamp [21] 100 28.87 0.9019 0.9990 0.9976 0.6646
SepMark [23] 30 35.73 0.9876 0.9957 0.9643 0.5086
TrustMark [2] 100 41.19 0.9922 0.9996 0.9955 0.7868
EditGuard-C [28] 64 37.27 0.9332 0.9991 0.5482 0.4925
EditGuard-D [28] 64 32.30 0.8199 0.9999 0.5444 0.4975
Robust-Wide [7] 64 41.58 0.9923 1.0000 0.9944 0.4951
VINE [13] 100 36.04 0.9874 0.9997 0.9986 0.5012
@@
WAM [19] 32 39.32 0.9791 1.0000 0.9986 0.8979
MaskMark-D (Ours) 32 39.55 0.9814 1.0000 1.0000 0.9998
MaskMark-ED (Ours) 32 39.52 0.9828 1.0000 1.0000 1.0000
@@
The local watermarking results for all methods are presented in Figure 3. Global watermarking methods suffer a significant drop in extraction accuracy as the watermarked region shrinks, revealing their weakness in local watermarking tasks. In contrast, both WAM and MaskMark maintain high accuracy even with small watermarked areas, and MaskMark consistently outperforms WAM, especially under distortions, demonstrating greater robustness. Between our variants, MaskMark-ED performs better when the watermarked area is small, with the performance gap narrowing as the watermarked area increases.
Figure 3: Watermark extraction performance of different methods under different ratios of watermarked areas. The intervals of ratios are: 1-5%, 5-10%, …, 95-99%, 99-100%. We select the average value for each interval’s ratios to stand for the interval (e.g., 3% for 1-5%).
Visualized Results.
The visualized watermark patterns embedded by MaskMark-D and MaskMark-ED are shown in Figure 15 in Appendix D.6.
@@ 4.3 Watermark Localization Comparison
We further compare MaskMark with EditGuard [28] and WAM [19], two methods with watermark localization capabilities. Localization performance is evaluated using the local watermarking dataset described in Sec. 4.2, with Intersection-over-Union (IoU) metrics computed between the predicted and ground-truth regions for both the watermarked and unwatermarked areas.
@@ Results.
The localization results of different methods are presented in Figure 4. First, MaskMark consistently achieves the best localization performance across nearly all watermark ratios and distortion conditions, with WAM showing a noticeable performance gap and EditGuard performing significantly worse. Second, MaskMark-ED outperforms MaskMark-D in localizing small regions, whether watermarked or unwatermarked, especially under distortion conditions. Interestingly, EditGuard-C shows an unusual rise in localization performance for unwatermarked regions as the watermark area increases under non-distorted conditions, a behavior not observed in EditGuard-D, possibly due to overfitting from being trained solely on clean data. The visualized localization results of different methods are provided in Appendix D.7.
Figure 4: Localization performance of different methods under different ratios of watermarked areas.
4.4 Performance Comparisons of Embedding Multiple Watermarks
@@
We further compare MaskMark-ED with WAM [19] under a multi-watermark setting by embedding up to five distinct watermarks into separate masked regions. The evaluation considers extraction accuracy and localization performance under a stricter masking constraint (5% area per region). We report the average watermark extraction accuracy across all embedded watermarks and the mean IoU of the predicted watermark regions. Detailed setup is provided in Appendix C.2.4.
@@ Results.
Figure 5 presents the comparison results. First, our MaskMark-ED consistently outperforms WAM in both watermark extraction and localization, across all tested numbers of embedded watermarks and under various distortion conditions. Second, despite being trained solely with single-watermark supervision, MaskMark-ED generalizes well to multi-watermark settings, demonstrating strong scalability. However, under geometric distortions, MaskMark-ED’s extraction accuracy degrades as the number of embedded watermarks increases. This is mainly due to spatial transformations (particularly rotation) shrinking the watermark regions in the image corners, reducing the effective area available for extraction (see Figure 20 in Appendix D.8).
@@ 4.5 More Analysis
While WAM is constrained to 32-bit watermarks, our MaskMark readily scales to longer bit lengths. Detailed results are provided in Figure 12 in Appendix D.3. MaskMark maintains high extraction accuracy across various bit lengths. Even at 64 bits, the accuracy experiences only a slight drop and still outperforms WAM at 32 bits. While the accuracy decline becomes more noticeable at 128 bits, MaskMark consistently surpasses WAM under both no distortion and valuemetric distortion conditions. The only exception arises under geometric distortion, where a performance gap emerges when the watermarked region covers between 5% and 75%. These results underscore the scalability and optimization-friendly design of our method, emphasizing its practical advantages.
@@
While MaskMark is trained on a broad range of common distortions, it is impractical to cover all possible scenarios. Fortunately, our framework supports task-specific fine-tuning, allowing it to adapt to new and emerging threats. As a demonstration, we consider adaptive attacks based on variational autoencoders (VAEs) [10], which reconstruct images and may inadvertently remove watermark signals. To counter this, we expand the distortion pool during fine-tuning by incorporating VAE modules from Stable Diffusion v1-4 [16], Bmshj18 [1], and Cheng20 [3]. Training details are provided in Appendix C.1.1. Fine-tuning for 20k steps on a single A6000 GPU ($\sim$5 hours) significantly improves robustness against these attacks, as shown in Figure 13 in Appendix D.4, outperforming existing baselines. While there is a slight drop in robustness to other distortions, MaskMark still maintains strong overall performance. These results demonstrate the adaptability of our approach to specific threats.
@@
We evaluate watermark extraction performance using three masking strategies when processing the watermarked image for extraction: (1) a full mask, which uses the entire image; (2) a predicted mask, which focuses on regions predicted to contain the watermark; and (3) a ground-truth mask, which focuses on the actual watermark-embedded regions. As shown in Figure 14 in Appendix D.5, when watermarks are embedded in small regions, using the full mask significantly reduces accuracy. In contrast, the predicted and ground-truth masks progressively improve performance. This highlights the importance of localizing watermarked regions before extraction to reduce interference from irrelevant content.
@@
We compare training and inference costs of MaskMark with EditGuard [28] and WAM [19], both supporting watermark extraction and localization, summarized in Table 2. MaskMark has fewer parameters than WAM, with similar encoder and decoder sizes; WAM allocates most parameters to the decoder due to its more complex task. In training, MaskMark requires only about 1/15 of WAM’s TFLOPs. EditGuard doesn’t report training cost, but its data consumption (steps×$\times$×batch) already reaches 60% of MaskMark ’s, despite using fewer strong distortions for robustness. Fewer parameters don’t always mean lower inference overhead, as different image processing methods affect memory and speed. Although EditGuard has the fewest parameters, it uses more memory and runs slower due to embedding both localization and copyright watermarks. In contrast, MaskMark and WAM have lower latency with similar memory use.
@@
EditGuard [28] 3.6 2.6 6.2 250K 4 - - - - 0.074 s 0.080 s 0.154 s 2.15 GB
WAM [19] 1.1 96.0 97.1 3,680K 16 8 V100 90 h $1.13\times 10^{16}$ 0.015 s 0.017 s 0.032 s 2.33 GB
MaskMark (Ours) 31.1 32.2 63.3 100K 16 1 A6000 20 h $7.74\times 10^{14}$ 25.84 GB 0.009 s 0.022 s 0.031 s 2.21 GB